OpenStudy (anonymous):

(1 pt) Evaluate the integral \int\!\!\!\int_R (x^{2}-2y^{2}) \, dA, where R is the first quadrant region between the circles of radius 4 and radius 5.

OpenStudy (anonymous):

Use polar coords?\[ x=r\cos(\theta)\\ y=r\sin(\theta)\\ dA = rdrd\theta\\ R= [4,5]\times [0,2\pi] \]